# Steering Large Language Models with Feature Guided Activation Additions

### v2

@@ Appendix D Cosine similarity of steering vectors
Appendix E Gemma-2-9B Steering Results
Figure E.1: Plots showing mean BCS with 95% confidence intervals for the CAA, SAE, SAE-TS and FGAA steering methods on 9 tasks, for Gemma-2-9B.
Appendix F Trade-off Curves
Figure F.1: Score trade-off curves for Gemma-2-2B, plotting both Coherence and Behavioral scores against increasing steering scale values. Each line tracks a distinct steering technique, with the optimal results appearing in the upper-right quadrant, where both Coherence and Behavioral metrics reach their highest values.
Figure F.2: Score trade-off curves for Gemma-2-9B, plotting both Coherence and Behavioral scores against increasing steering scale values. Each line tracks a distinct steering technique, with the optimal results appearing in the upper-right quadrant, where both Coherence and Behavioral metrics reach their highest values.
Appendix G Normalisation of $v_{target}$
@@
Appendix H Family of BOS features
@@
Table H.1: Identified BOS Features from Gemma-2-2B 16k SAE (non-exhaustive). Descriptions marked with an asterisk (*) are the authors’ interpretations. Uninterpretable features are not included.

### v3

@@ Appendix D Cosine similarity of steering vectors
Appendix E Trade-off Curves
Figure E.1: Score trade-off curves for Gemma-2-2B, plotting both Coherence and Behavioral scores against increasing steering scale values. Each line tracks a distinct steering technique, with the optimal results appearing in the upper-right quadrant, where both Coherence and Behavioral metrics reach their highest values.
Figure E.2: Score trade-off curves for Gemma-2-9B, plotting both Coherence and Behavioral scores against increasing steering scale values. Each line tracks a distinct steering technique, with the optimal results appearing in the upper-right quadrant, where both Coherence and Behavioral metrics reach their highest values.
Appendix F Normalisation of $v_{target}$
@@
Appendix G Family of BOS features
Table G.1: Identified BOS Features from Gemma-2-2B 16k SAE (non-exhaustive). Descriptions marked with an asterisk (*) are the authors’ interpretations. Uninterpretable features are not included.